# LLäMmlein [Uncaptioned image]: Transparent, Compact and Competitive German-Only Language Models from Scratch

### v2

@@ Abstract
We transparently create two German-only decoder models, LLäMmlein 120M and 1B, from scratch and publish them, along with the training data, for the (German) NLP research community to use111 https://www.informatik.uni-wuerzburg.de/datascience/projects/nlp/llammlein/ . The model training involved several key steps, including data preprocessing/filtering, the creation of a German tokenizer, the training itself, as well as the evaluation of the final models on various benchmarks, also against existing models. Throughout the training process, multiple checkpoints were saved in equal intervals and analyzed using the German SuperGLEBer benchmark to gain insights into the models’ learning process.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success, yet this progress is predominantly centered on English. Other languages, including German, lag behind due to limited competition, reduced investment, and a lack of transparency in training data, code, and detailed results (also see Pfister and Hotho, 2024). While smaller German-only models do exist, such as BERTs or smaller GPTs Chan et al. (2020); Scheible et al. (2020), or the contemporaneous effort by Idahl (2024), many are closed or undocumented, and few robust, openly accessible German LLMs have been built from scratch with full transparency. Most German-capable LLMs are either multilingual models (e.g., mGPT (2024)) or English models adapted to German (e.g., LeoLM (2023b), BübleLM (2024), or bloom-clp (2023)). Others, like Mistral (2023), share little about their (German) training data, making it difficult to understand which resources and pretraining strategies are most suitable for developing strong German LLMs from scratch. This lack of traceability hampers the community’s ability to identify, curate, and refine German data, and to examine how corpus and training choices affect model quality. Subtle issues - such as performance deteriorations on non-English downstream tasks Virtanen et al. (2019) or poor handling of German’s complex grammar and morphology (Mielke et al. (2019), also Appendix A) - remain common, even in state-of-the-art models like Llama 3 Dubey et al. (2024), which can revert to English despite a German context, or sometimes sound like machine-translated from English [^2] https://www.reddit.com/r/LocalLLaMA/comments/1bfce18/still_didnt_found_a_better_small_german_llm_anyone/.
@@
To achieve this, we: (1) clean and filter a large German dataset derived from RedPajama V2 Computer (2023), ensuring high-quality input, (2) construct a dedicated German tokenizer (32k tokens) fitted on varying data amounts to compare against existing German tokenizers, (3) pretrain two exclusively German autoregressive LLMs (120M and 1B) and release incremental checkpoints, inspired by Biderman et al. (2023), to inform efficient stopping criteria and shed light on learning dynamics, and (4) evaluate the models on a range of tasks (SuperGLEBer Pfister and Hotho (2024), lm-evaluation-harness-de Plüster (2023a); Gao et al. (2021)) to benchmark performance against existing models.
@@ 2 Methodology
Pretraining and evaluating a German LLM from scratch, end-to-end involves several steps, including dataset preprocessing (section 2.1.2), tokenizer fitting (section 3.1), model pretraining (section 2.2), model evaluation using a comprehensive German benchmark, as well as the standard translated prompt-based few-shot QA tasks (section 2.3), and exemplary downstream adaptations (section 2.4).
@@ 2.1 Dataset
RedPajama V2 is an open [^3] https://commoncrawl.org/terms-of-use, multilingual dataset designed for training large language models Computer (2023). It consists of over 100 billion text documents collected from 84 CommonCrawl snapshots between 2014 and 2023 and encompasses multiple languages, including English, German, French, Italian, and Spanish. The dataset was originally preprocessed using the CCNet pipeline Wenzek et al. (2020) leading to about 30 billion overall documents further enriched with quality signals and duplicate indicators. Using perplexity of a language model, the RedPajama dataset was divided into three quality categories, in descending order of quality: head, middle, and tail. Following a manual inspection of a randomly selected subset, the head and middle partitions were deemed to contain sufficiently high-quality German texts suitable for continued use. In contrast, the tail partition exhibited inconsistent quality and was consequently excluded from further training.
@@ 2.1.1 Dataset Analysis
The aim of the following preliminary dataset analysis is to gain a deeper understanding of the German portion of the dataset used. The “official” estimate of the size of the German segment within the RedPajamaV2 dataset, derived through extrapolation from a smaller sample analyzed with Mistral-7B, is approximately 3 trillion German tokens Computer (2023). We will perform an exploratory analysis of the dataset to gain a clearer understanding of the actual amount of German data it contains, alongside its domain distribution and the most prevalent data sources.
@@
Figure 1 breaks down the distribution of each partition, i.e. head unique, middle unique, head duplicate and middle duplicate separately. The first occurence of a sample is considered unique, while all subsequent appearances are marked as duplicates. The middle unique partition contains the largest amount of data, with approximately 1.2 billion samples, which corresponds to 45% of the full dataset. The head unique partition, by comparison, includes around 400 million fewer samples.
Overall, most samples are unique, as in appearing only once (1.9 billion samples) and only significantly less are marked as duplicates, appearing a second or more times (777 million samples) across the entire dataset. The most common token-per-document count can be found at nine, with approximately 3.6 million occurrences across the entire dataset. A second peak (most prominent for the unique split) occurs at around 100 tokens per document. In total, the 2.7 trillion German tokens are distributed across samples with lengths ranging from 1 to 1 034 79910347991\,034\,7991 034 799 tokens, averaging approximately $1000$ tokens per sample.
@@
Wikipedia clearly stands out as the largest contributor, with a combined total of over 11.5 million samples. Among these, about 10 million entries belong to the head category, while about 1.45 million stem from the middle partition. This distribution aligns with the fact that the split into head, middle and tail was created using a perplexity score criterion based on a language model trained on Wikipedia Computer (2023) - consequently, texts closer in style to Wikipedia tend to be ranked higher. Besides Wikipedia, it is evident that news websites also constitute a significant portion of the dataset. For the middle split, “welt.de” emerges as the most frequent domain, contributing around 2.47 million samples. With the exception of domains like eBay, hotels and Perlentaucher, the list is largely dominated by general news outlets.
@@ 2.1.2 Further Dataset Preprocessing
To remove common web boilerplate, such as GDPR notices or similar repetitive content, we utilize a paragraph-level deduplication scheme powered by Dolma - a framework that enables efficient deduplication through a Rust-based Bloom filter Soldaini et al. (2024). This ensures that highly redundant text is filtered out, improving the overall quality and diversity of the dataset. This approach may inadvertently over-remove valid and relevant content, such as short texts mistakenly treated as entire paragraphs being removed across the dataset. To mitigate this, and to preserve meaningful short text sections - such as lists or frequently occurring itemized phrases that are contextually significant - we excluded texts containing fewer than three words from the deduplication process.
Despite these efforts, we found that some unusual artifacts, such as long sequences of guitar chords, remained, as they scored low perplexity (41.2) compared to the average perplexity of 206.35 in the respective snapshot (2014_52) and were therefore not removed by the preliminary quality filter. To address this, we built a token-to-word ratio filter. Here, we compared the word count (whitespace separated) with the token count using the German GPT-2 tokenizer Schweter (2020). According to our intuition, an usual ratio between the two counts indicates abnormal or low-quality text, whereas a close match suggests valid German content. A simple example illustrates this clearly: The phrase “Der Himmel ist blau” consists of 4 words and 4 tokens, so it is not removed by our filter. In contrast, “/de/c/trebic-unesco” counts as 1 word but 11 tokens, and should therefore be excluded by this token-to-word ratio filter.
Preliminary examinations and manual review suggested a ratio of tokens to words of eight as a valid threshold. Thus, paragraphs with ratios exceeding this threshold were excluded from the dataset.
@@ 2.2 Model Pretraining Framework
While there are several existing resources and repositories for training a LLM from scratch444a small subset: https://github.com/Hannibal046/Awesome-LLM#llm-training-frameworks, we chose the TinyLlama GitHub repository as the backbone of our project Zhang et al. (2024). It was used to pretrain the 1B English TinyLlama model from scratch before and builds upon the lit-gpt repository AI (2023), which provides robust tooling for data preparation, fine-tuning, pretraining, and deploying LLMs using PyTorch Lightning.
It includes all relevant features such as multi-GPU and multi-node distributed training with FSDP as well as FlashAttention-2 Dao (2024). Additionally, it provides scripts to convert the models into HuggingFace format for easier use and distribution.
@@ 2.3.1 Intermediate Checkpoint Evaluation
To get a better understanding of the training, we monitor the progress and regularly evaluate intermediate checkpoints on six representative SuperGLEBer tasks Pfister and Hotho (2024). These tasks were selected to encompass a range of problem types to assess our model’s performance. Within classification, (1) Natural Language Inference(NLI) Conneau et al. (2018) requires determining whether a hypothesis is entailed, neutral, or contradictory to a premise; (2) FactClaiming Comments Risch et al. (2021) involves binary classification of fact-checkable claims; (3) DB Aspect Wojatzki et al. (2017) addresses multi-label categorization and polarity detection in input sentences; and (4) WebCAGe Henrich et al. (2012) tests if a given word’s sense aligns across two contexts. For sequence tagging, (5) EuroParl Faruqui and Padó (2010) evaluates token-level classification in European Parliament data. Finally, in the sentence similarity domain, (6) PAWSX Liang et al. (2020) challenges the model to detect paraphrases via vector representations.
@@ 2.3.2 Final Model Evaluation
To assess general knowledge and abilities, we evaluated our final models on the full SuperGLEBer benchmark (29 tasks across classification, sequence tagging, question answering, and sentence similarity) Pfister and Hotho (2024), as well as machine-translated, prompt-based, few-shot QA tasks using the lm-evaluation-harness-de by Plüster (2023a) (if not stated otherwise, they are evaluated measuring unnormalized and Byte-length normalized accuracy Gao et al. (2021)): (1) ARC-Challenge-DE Clark et al. (2018): Grade-school science questions ($1471$ samples). Few-shot evaluation with 25 samples. (2) MMLU-DE Hendrycks et al. (2021): $6829$ multiple-choice questions across 57 topics (e.g., math, medicine, law). Few-shot evaluation with five samples. (3) HellaSwag-DE Zellers et al. (2019): Commonsense reasoning dataset with 11 0001100011\,00011 000 translated samples, featuring incomplete sentences with multiple-choice completions. Few-shot evaluation with ten samples. (4) TruthfulQA-DE Lin et al. (2022): 817 questions across 38 categories (e.g., health, law) designed to evaluate model truthfulness, particularly in handling misconceptions. Zero-shot evaluation. Performance here is measured using MC1 (Single-true): accuracy in selecting a single correct answer and MC2 (Multi-true): the normalized probability assigned to all correct answers Gao et al. (2021).
@@ 3.1 Tokenizer
We follow the TinyLlama setup and fit a Llama 2 Byte-Pair Encoding (BPE) tokenizer with a 32 0003200032\,00032 000-token vocabulary. We trained three tokenizers on different amounts of data:
@@ 3.2.2 LLäMmlein 1B
LLäMmlein 1B was trained on the filtered, head and middle unique partitions, resulting in a dataset of 1.3T unique and 3T total tokens. The training setup featured a maximum learning rate of 6e-4, a global batch size of 1024 (per device batch size of 16), and was executed on 64 A100 80GB GPUs across 8 nodes over 32 days (50 0005000050\,00050 000 GPU hours).
@@ 4.2.2 LLäMmlein 1B
We compared LLäMmlein 1B’s performance on the SuperGLEBer benchmark to the best-performing models for each task and similarly sized models (Table 3). While not always securing the top spot, it remains competitive across tasks, even against much larger models. As with the 120M model, LLäMmlein 1B trails in sentence similarity tasks like PAWSX. However, it achieves competitive results for EuroParl. Examining task progress over time reveals noticeable improvements across all tasks, except for FactClaiming. Compared to the LLäMmlein 120M model, Spearman correlation analysis indicated significant positive relationships between training time and performance for all remaining tasks. In particular, also for EuroParl ($r$ = 0.431, $p$ = 0.009) and WebCAGe ($r$ = 0.92, $p$ < 0.0001), suggesting that LLäMmlein 1B continues to benefit from extended training. Nevertheless, across all SuperGLEBer tasks, the advantage of extended pretraining diminished after roughly 30% of the pretraining data was processed. From this state, despite a slow decline in loss, no significant improvements were observed across the 29 downstream tasks (Figure 6). To investigate further, we evaluated the checkpoint where SuperGLEBer performance plateaued, along with its instruction-tuned variants on generative tasks Plüster (2023a). Interestingly, while SuperGLEBer performance stagnated, generative benchmark results (Table 9) continued to improve on average, likely due to enhanced autoregressive language modeling capabilities.

### v3

@@ Abstract
We transparently create two German-only decoder models, LLäMmlein 120M and 1B [^1] After submission we also released a new 7B model, from scratch and publish them, along with training data, for the (German) NLP research community to use222 https://www.informatik.uni-wuerzburg.de/datascience/projects/nlp/llammlein/ . The model training involved several key steps, including data preprocessing/filtering, the creation of a German tokenizer, the training itself, as well as the evaluation of the final models on various benchmarks, also against existing models. Throughout the training process, multiple checkpoints were saved in equal intervals and analyzed using the German SuperGLEBer benchmark to gain insights into the models’ learning process.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success, yet this progress is predominantly centered on English. Other languages, including German, lag behind due to limited competition, reduced investment, and a lack of transparency in training data, code, and detailed results (also see Pfister and Hotho, 2024). While smaller German-only models do exist, such as BERTs or smaller GPTs Chan et al. (2020); Scheible et al. (2020), or the contemporaneous effort by Idahl (2024), many are closed or undocumented, and few robust, openly accessible German LLMs have been built from scratch with full transparency. Most German-capable LLMs are either multilingual models (e.g., mGPT Shliazhko et al. (2024)) or English models adapted to German (e.g., LeoLM Plüster (2023b), BübleLM Delobelle et al. (2024), or bloom-clp Ostendorff and Rehm (2023)). Others, like Mistral Jiang et al. (2023), share little about their (German) training data, making it difficult to understand which resources and pretraining strategies are most suitable for developing strong German LLMs from scratch. This lack of traceability hampers the community’s ability to identify, curate, and refine German data, and to examine how corpus and training choices affect model quality. Subtle issues – such as performance deteriorations on non-English downstream tasks Virtanen et al. (2019) or poor handling of German’s complex grammar and morphology (Mielke et al. (2019), also Appendix A) – remain common, even in state-of-the-art models like Llama 3 Dubey et al. (2024), which can revert to English despite a German context, or sometimes sound like machine-translated from English [^3] https://www.reddit.com/r/LocalLLaMA/comments/1bfce18/still_didnt_found_a_better_small_german_llm_anyone/.
@@
To achieve this, we: (1) clean and filter a large German dataset derived from RedPajama V2 Weber et al. (2024), ensuring high-quality input, (2) construct a dedicated German tokenizer (32k tokens) fitted on varying data amounts to compare against existing German tokenizers, (3) pretrain two exclusively German autoregressive LLMs (120M and 1B) and release incremental checkpoints, inspired by Biderman et al. (2023), to inform efficient stopping criteria and shed light on learning dynamics, and (4) evaluate the models on a range of tasks (SuperGLEBer Pfister and Hotho (2024), lm-evaluation-harness-de Plüster (2023a); Gao et al. (2021)) to benchmark performance against existing models.
@@ 2 Methodology
Pretraining and evaluating a German LLM from scratch, end-to-end involves several steps, including dataset preprocessing (section 2.1.2), tokenizer fitting (section 3.1), model pretraining (section 2.2), model evaluation using a comprehensive German benchmark, as well as multiple translated prompt-based few-shot QA tasks (section 2.3), and exemplary downstream adaptations (section 2.4).
@@ 2.1 Dataset
RedPajama V2 is an open [^4] https://commoncrawl.org/terms-of-use, multilingual dataset designed for training large language models Weber et al. (2024). It consists of over 100 billion text documents collected from 84 CommonCrawl snapshots between 2014 and 2023 and encompasses multiple languages, including English, German, French, Italian, and Spanish. The dataset was originally preprocessed using the CCNet pipeline Wenzek et al. (2020) leading to about 30 billion overall documents further enriched with quality signals and duplicate indicators. Using perplexity of a language model, the RedPajama dataset was divided into three quality categories, in descending order of quality: head, middle, and tail. Following a manual inspection of a randomly selected subset, the head and middle partitions were deemed to contain sufficiently high-quality German texts suitable for continued use. In contrast, the tail partition exhibited inconsistent quality and was consequently excluded from further training.
@@ 2.1.1 Dataset Analysis
The aim of the following preliminary dataset analysis is to gain a deeper understanding of the German portion of the dataset used. The “official” estimate of the size of the German segment within the RedPajamaV2 dataset, derived through extrapolation from a smaller sample analyzed with Mistral-7B, is approximately 3 trillion German tokens Weber et al. (2024). Following, we first perform an exploratory analysis of the dataset to gain a clearer understanding of the actual amount of German data it contains, alongside its domain distribution and the most prevalent data sources.
@@
Figure 1 breaks down the distribution of each partition, i.e. head unique, middle unique, head duplicate and middle duplicate separately. The first occurence of a document is considered unique, while all subsequent appearances are marked as duplicates. The middle unique partition contains the largest amount of data, with approximately 1.2 billion samples, which corresponds to 45% of the full dataset. The head unique partition, by comparison, includes around 400 million fewer samples.
Overall, most samples are unique (1.9 billion samples) and only significantly less are marked as duplicates, appearing a second or more times (777 million samples) across the entire dataset. The most common token-per-document count can be found at nine, with approximately 3.6 million occurrences in the dataset. A second peak (most prominent for the unique split) occurs at around 100 tokens per document. In total, the 2.7 trillion German tokens are distributed across samples with lengths ranging from 1 to 1 034 79910347991\,034\,7991 034 799 tokens, averaging approximately $1000$ tokens per sample.
@@
Wikipedia clearly stands out as the largest contributor, with a combined total of over 11.5 million samples. Among these, about 10 million entries belong to the head category, while about 1.45 million stem from the middle partition. This distribution aligns with the fact that the split into head, middle and tail was created using a perplexity score criterion based on a language model trained on Wikipedia Weber et al. (2024) – consequently, texts closer in style to Wikipedia tend to be ranked higher. Besides Wikipedia, it is evident that news websites also constitute a significant portion of the dataset. For the middle split, “welt.de” emerges as the most frequent domain, contributing around 2.47 million samples. With the exception of domains like eBay, hotels and Perlentaucher, the list is largely dominated by general news outlets.
@@ 2.1.2 Further Dataset Preprocessing
To remove common web boilerplate, such as EU-specific “General Data Protection Regulation” (GDPR) notices or similar repetitive content, we utilize a paragraph-level deduplication scheme powered by Dolma – a framework that enables efficient deduplication through a Rust-based Bloom filter Soldaini et al. (2024). A Bloom filter is a probabilistic method that works similarly to a hash table, determining whether an element has already been previously encountered or not. This ensures that highly redundant text is filtered out, improving the overall quality and diversity of the dataset. This approach may inadvertently over-remove valid and relevant content, such as short texts mistakenly treated as entire paragraphs being removed across the dataset. To mitigate this, and to preserve meaningful short text sections – such as lists or frequently occurring itemized phrases that are contextually significant – we excluded paragraphs containing fewer than three words from the deduplication process.
Despite these efforts, we found that some unusual artifacts, such as long sequences of guitar chords, remained, as they scored low perplexity (41.2) compared to the average perplexity of 206.35 in the respective snapshot (2014_52) and were therefore not removed by the preliminary quality filter. To address this, we built a token-to-word ratio filter. Here, we compared the word count (whitespace separated) with the token count using the German GPT-2 tokenizer Schweter (2020). According to our intuition, a usual ratio between the two counts indicates abnormal or low-quality text, whereas a close match suggests valid German content. A simple example illustrates this clearly: The phrase “Der Himmel ist blau” consists of 4 words and 4 tokens, so it is not removed by our filter. In contrast, “/de/c/trebic-unesco” counts as 1 word but 11 tokens, and should therefore be excluded by this token-to-word ratio filter.
Preliminary examinations and manual review suggested a ratio of tokens to words of eight as a valid threshold. Thus, paragraphs exceeding this threshold were excluded from the dataset.
@@ 2.2 Model Pretraining Framework
While there are several existing resources and repositories for training an LLM from scratch555a small subset: https://github.com/Hannibal046/Awesome-LLM#llm-training-frameworks, we chose the TinyLlama GitHub repository as the backbone of our project Zhang et al. (2024). It was used to pretrain the 1B English Llama 2-based Touvron et al. (2023) TinyLlama model from scratch before and builds upon the lit-gpt repository AI (2023), which provides robust tooling for data preparation, fine-tuning, pretraining, and deploying LLMs using PyTorch Lightning.
It includes all key features such as multi-GPU and multi-node distributed training with FSDP as well as FlashAttention-2 Dao (2024). In addition, it provides scripts to convert the models into HuggingFace format for easier use and distribution.
We modified the codebase [^6] https://github.com/LSX-UniWue/LLaMmlein for our requirements: 1. We significantly improved the data loader speed by adding various layers of caching. 2. We enable training directly from a directory of jsonl-files, without any prior preprocessing. 3. Most importantly, inspired by our reviewers’ feedback, we retrained our models logging exact datapoints as they enter the model. This allows us to correlate the training data and its order for each of our published (intermediate) checkpoints [^7] Since the performance of the original model (without data logging) and the new model (with data logging) did not differ significantly, we kept the original scores throughout the paper..
@@ 2.3.1 Intermediate Checkpoint Evaluation
To get a better understanding of the training, we monitor the progress and regularly evaluate intermediate checkpoints on six representative SuperGLEBer tasks Pfister and Hotho (2024) following a finetuning to the task from the respective checkpoint. These tasks were selected to encompass a range of problem types to assess our model’s performance. Within classification, (1) Natural Language Inference(NLI) Conneau et al. (2018) requires determining whether a hypothesis is entailed, neutral, or contradictory to a premise; (2) FactClaiming Comments Risch et al. (2021) involves binary classification of fact-checkable claims; (3) DB Aspect Wojatzki et al. (2017) addresses multi-label categorization and polarity detection in input sentences; and (4) WebCAGe Henrich et al. (2012) tests if a given word’s sense aligns across two contexts. For sequence tagging, (5) EuroParl Faruqui and Padó (2010) evaluates Named Entity Recognition on European Parliament data. Finally, in the sentence similarity domain, (6) PAWSX Liang et al. (2020) challenges the model to detect paraphrases via vector representations.
@@ 2.3.2 Final Model Evaluation
To assess general knowledge and abilities, we evaluated our final models on the full SuperGLEBer benchmark (29 tasks across classification, sequence tagging, question answering, and sentence similarity) Pfister and Hotho (2024), as well as machine-translated, prompt-based, few-shot QA tasks using the lm-evaluation-harness-de by Plüster (2023a) (if not stated otherwise, they are evaluated measuring unnormalized and Byte-length normalized accuracy Gao et al. (2021)): (1) ARC-Challenge-DE Clark et al. (2018): Grade-school science questions ($1471$ samples). Few-shot evaluation with 25 samples. (2) MMLU-DE Hendrycks et al. (2021): $6829$ multiple-choice questions across 57 topics (e.g., math, medicine, law). Few-shot evaluation with five samples. (3) HellaSwag-DE Zellers et al. (2019): Commonsense reasoning dataset with 11 0001100011\,00011 000 translated samples, featuring incomplete sentences with multiple-choice completions. Few-shot evaluation with ten samples. (4) TruthfulQA-DE Lin et al. (2022): 817 questions across 38 categories designed to evaluate model truthfulness, particularly in handling misconceptions in a zero-shot evaluation setting. Performance here is measured using MC1 (Single-true): accuracy in selecting a single correct answer and MC2 (Multi-true): the normalized probability assigned to all correct answers Gao et al. (2021).
@@ 3.1 Tokenizer
We follow the TinyLlama setup and fit an Llama 2 Byte-Pair Encoding (BPE) tokenizer with a 32 0003200032\,00032 000-token vocabulary Touvron et al. (2023). We trained three tokenizers on different amounts of data:
@@ 3.2.2 LLäMmlein 1B
LLäMmlein 1B was trained on the filtered, head and middle unique partitions, resulting in a dataset of 1.3T unique tokens, and 3T overall tokens seen during training. The training setup featured a maximum learning rate of 6e-4, a global batch size of 1024 (per device batch size of 16), and was executed on 64 A100 80GB GPUs across 8 nodes over 32 days (50 0005000050\,00050 000 GPU hours).
@@ 4.1 Tokenizer
To validate this finding, we repeated the experiment using a variety of disjoint training datasets and different test sets, including snapshots from different time periods as well as random internet texts. Interestingly, we consistently observed the same outcome across all variations.
@@ 4.2.2 LLäMmlein 1B
We compared LLäMmlein 1B’s performance on the SuperGLEBer benchmark to the best-performing models for each task and similarly sized models (Table 3). All models and checkpoints are evaluated after a task-specific finetuning, following SuperGLEBer evaluation protocol. While not always securing the top spot, it remains competitive across tasks, even against much larger models. As with the 120M model, LLäMmlein 1B trails in sentence similarity tasks like PAWSX. However, it achieves competitive results for EuroParl. Examining task progress over time reveals noticeable improvements across all tasks, except for FactClaiming. Compared to the LLäMmlein 120M model, Spearman correlation analysis indicated significant positive relationships between training time and performance for all remaining tasks. In particular, also for EuroParl ($r$ = 0.431, $p$ = 0.009) and WebCAGe ($r$ = 0.92, $p$ < 0.0001), suggesting that LLäMmlein 1B continues to benefit from extended training. Nevertheless, across all SuperGLEBer tasks, the advantage of extended pretraining diminished after roughly 30% of the pretraining data was processed. From this state, despite a slow decline in loss, no significant improvements were observed across the 29 downstream tasks (Figure 6). To investigate further, we evaluated the checkpoint where SuperGLEBer performance plateaued, along with its instruction-tuned variants on generative tasks Plüster (2023a). Interestingly, while SuperGLEBer performance stagnated, generative benchmark results (Table 9) continued to improve on average, likely due to enhanced autoregressive language modeling capabilities.